OpenStudy (ashleynguyenx3):

A spherical balloon is being filled in such a way that the surface area is increasing at a rate of 20 cm^2/sec when the radius is 2 meters. At what rate is air being pumped in the balloon when the radius is 2 meters?

OpenStudy (anonymous):

Surface area of a sphere?

OpenStudy (ashleynguyenx3):

A=4pi(r^2)

OpenStudy (anonymous):

Good good. And we know \(\large \frac{dA}{dt}\) = 20 cm^2/sec

OpenStudy (ashleynguyenx3):

So would I find the derivative of the area formula?

OpenStudy (anonymous):

Take the derivative of A = 4pi(r^2) to get \(\large \frac{dA}{dr}\)

OpenStudy (anonymous):

Derivative with respect to r, I mean.

OpenStudy (ashleynguyenx3):

dA/dr = 4pi(2r)?

OpenStudy (anonymous):

Right, and 4pi(2r) = 8pi*r \(\Large \frac{dA}{dt} /div \frac{dA}{dr} = \frac{dr}[dt}\)

OpenStudy (ashleynguyenx3):

Wait, dA/dt=8pi(r)dr/dt?

OpenStudy (ashleynguyenx3):

So would I just then plug in the r and dA/dt to get my answer?

OpenStudy (anonymous):

Okay, Ashley. You start out knowing dA/dt. You are asked to find out dV/dt. Take derivative of A=4pir^2 to get dA/dr dA/dt / divided by dA/dr gives you dr/dt Now look at your volume equation: V = (4/3)pi*r^3 Take the derivative with respect to r to get dV/dr dr/dt * dV/dr = dV/dt

OpenStudy (anonymous):

How much of that makes sense to you?

OpenStudy (ashleynguyenx3):

dV/dr=4pi(r^2)?

OpenStudy (ashleynguyenx3):

I'm so confused.

OpenStudy (anonymous):

You are correct =) dV/dr = 4*pi*r^2

OpenStudy (anonymous):

I just want to chime in here. I am doing related rates right now as well and SmoothMath, you made is very clear

OpenStudy (anonymous):

Ah, great =) Very good to hear.

OpenStudy (ashleynguyenx3):

So do I just multiply the two equations together to find dV/dt?

OpenStudy (anonymous):

For both of you, I would say, make sure you're clear about what you are starting with and what you are trying to get to. In this problem, we start with how quickly the surface area is changing, which is dA/dt And I am trying to figure out how much air is being blown into the balloon, which is how quickly the volume is changing, so dV/dt How do I get from dA/dt to dV/dt \(\Large \frac{dA}{dt}*\frac{dr}{dA}*\frac{dV}{dr}\)

OpenStudy (anonymous):

This is just like chemistry lol

OpenStudy (anonymous):

I agree =)

OpenStudy (anonymous):

Dimensional analysis.

OpenStudy (anonymous):

Ask me any questions you have please, Ashley. I want you to understand.

OpenStudy (ashleynguyenx3):

How would you get dr/dA?

OpenStudy (anonymous):

Okay, so I wrote that as \(\Large *\frac{dr}{dA}\) but you could also do \(\Large \div \frac{dA}{dr}\) and that would be the same thing. Do you know how to find dA/dr?

OpenStudy (ashleynguyenx3):

Didn't we find that earlier? dA/dr = 8pi(r)

OpenStudy (anonymous):

Right. Good.

OpenStudy (ashleynguyenx3):

So now what?

OpenStudy (anonymous):

What have you gotten so far?

OpenStudy (ashleynguyenx3):

dA/dr=8pi(r) dV/dr=4/3pi(r^2)

OpenStudy (anonymous):

And you also know, from the problem, that dA/dt = 20 cm^2/sec

OpenStudy (ashleynguyenx3):

Yes, and r=2

OpenStudy (anonymous):

Great. Soooo you can figure it out from there.

OpenStudy (anonymous):

\(\Large \frac{dA}{dt} \div \frac{dA}{dr}*\frac{dV}{dr}\)

OpenStudy (ashleynguyenx3):

So it'd just plug it in to get 20/16pi*16pi?

OpenStudy (anonymous):

I think your math may be getting mildly confused.

OpenStudy (anonymous):

\(\Large \frac{dA}{dt} \div \frac{dA}{dr} * \frac{dV}{dr}\) \(\Large \rightarrow 20 \div (8\pi*r) * (4\pi*r^2)\)

OpenStudy (ashleynguyenx3):

Oh, I plugged in r=2

OpenStudy (anonymous):

Great, that's fine.

OpenStudy (anonymous):

You got the wrong result though, so check your math.

OpenStudy (ashleynguyenx3):

I still got 20 / (16pi) (16pi)

OpenStudy (anonymous):

Oh okay. You mean \(\Large \frac{20}{16\pi}*16\pi\)

OpenStudy (anonymous):

That simplifies.

OpenStudy (ashleynguyenx3):

Yes, so I get 20 right? I kinda think we just solved dA/dt

OpenStudy (anonymous):

lol no don't worry. It's just a coincidence that it comes out the same.

OpenStudy (anonymous):

20 cm^3/sec is correct.

OpenStudy (ashleynguyenx3):

Okay, thanks. :D I honestly didn't get it.